# MonoDGP: Monocular 3D Object Detection with Decoupled-Query and Geometry-Error Priors

### v1

@@ 4.2 Main Results
As shown in Tab. 1, we have compared the most advanced monocular 3D object detection methods in recent years. Without any additional data, MonoDGP comprehensively surpasses the baseline MonoDETR [49], and achieves competitive results in most metrics. On test set and val set, compared to all existing methods, our performance for $AP_{3D}$ and $AP_{BEV}$ under three-level difficulties improves by -0.14%, +1.05%, +0.65%, +0.80%, +1.55%, +1.18%, +1.54%, +1.25%, +1.33%, +1.92%, +1.73%, +1.87%, respectively. Except that easy $AP_{BEV}$ on test set achieves the second-best lower than OccupancyM3D [29], other metrics all reach the SOTA performance.
It should be noted that we utilize early stopping to select the best checkpoint on val set, while we directly select the last few checkpoints and submit our results on test set. Considering the instability of the DETR-based model, on val set, we train MonoDGP 5 times and take the median of moderate $AP_{3D}$ scores as the final results. The rest of the ablation experiments are conducted in the same manner.
@@ 4.3 Ablation Study
Table 2: The design of region segmentation head. ‘w/o’ denotes enhancing visual and depth features without segment embeddings.
Segment Embeddings Threshold Val, $AP_{3D}$
@@
Region segmentation head.  For the multi-scale target region probabilities predicted by RSH, we set distinct thresholds to produce segment embeddings, which serve as input tokens for the depth encoder. The results are presented in Tab. 2. The introduction of segment embeddings strengths encoders’ ability to comprehend the contextual relationships among image pixels. Experimental findings further highlight the critical role of segment embeddings. Moreover, the threshold settings also influence 3D evaluation metrics. A threshold of 0.5 achieves superior results compared to thresholds of 0.3 and 0.7.
Table 3: Ablation study of different depth prediction modes on KITTI val set. The weighted fusion mode integrates the results from the depth map, direct depth, and geometric depth. "GD" denotes the geometric depth calculated by the formula $Z_{geo}=f*H_{3D}/h_{bbox}$.
Depth Predition Mode Val, $AP_{3D}$
@@
Geometric Depth 9.15 7.99 5.52
@@
Depth prediction mode.  Based on the same network architecture, we predict depth in different modes, and the ablation results are expressed in Tab. 3. The findings indicate that individual depth estimations, whether using direct depth, depth map, or geometric depth, generally perform poorly, especially for geometric depth alone. As only perspective transformation on the visual surface considered, the moderate $AP_{3D}$ for geometric depth is just 7.99% (all $AP_{3D}$s mentioned below default to the moderate level), which is significantly lower than other prediction modes. By fusing the direct depth, depth map and geometric depth, which is adopted by MonoDETR, the $AP_{3D}$ improves to 21.64%. When geometry errors are added to compensate the projection formula, it can be observed that the network’s performance achieves a substantial increase. The depth error, with an $AP_{3D}$ of 22.34%, wins the best performance among all depth prediction modes. The dimension height error, with 22.09% in $AP_{3D}$, achieves the second-best performance due to its skewed distribution. The bounding box height error only acquires 21.18% in $AP_{3D}$, obviously inferior to other geometry errors. This discrepancy can be attributed to the nonlinearity of the gradient backpropagation calculation.
Table 4: The main ablation study.
Mixup3D 2D Query Region Segment Geometric Depth Error Val, $AP_{3D}$
@@
Overall performance.  We explore the impact of each improvement technique on the baseline. The main ablation study is shown in Tab. 4. The MonoDGP employs four key techniques: mixup3D augmentation, decoupling and initialization of 2D object queries, region enhancement with segment embeddings, and depth error prediction based on geometric depth. Compared to the baseline, the $AP_{3D}$s under three-level difficulties gain an increase by +4.20%, +2.12% and +1.80%, respectively. The experiments validate the effectiveness of each technique utilized in MonoDGP.
@@ 5 Conclusion
In this paper, we propose a novel transformer-based monocular method called MonoDGP, which adopts geometry errors to correct the projection formula. We also comprehensively analyze the properties and effectiveness of geometry errors, from the perspectives of marginal feature extraction, data distribution, and gradient propagation. Our error prediction is expected to become a better alternative to multi-depth prediction, without complex implementation of various depth branches. Furthermore, MonoDGP introduces a 2D visual decoder for query initialization and constructs a region segmentation head for feature enhancement. These improvements have achieved superior performance on the KITTI benchmark. Additionally, we can extend our approach to dense depth map prediction for the target region, by only considering errors between geometric depth and surface points of the objects.

### v2

@@ 4.2 Main Results
As shown in Tab. 1, we have compared advanced monocular 3D detectors in recent years. Without additional data, MonoDGP comprehensively surpasses the baseline model MonoDETR [62] and achieves competitive results across most metrics. In particular, our method improves moderate $AP_{3D}$ by up to 1.55% on the test set and 1.73% on the validation set, respectively. Except that easy $AP_{BEV}$ on test set achieves the second-best lower than OccupancyM3D [39], which introduces extra data during the training phase, other metrics all reach the state-of-the-art performance.
It should be noted that we employ early stopping to select the best validation checkpoint, while we directly select the last few checkpoints for test submission. To mitigate the instability of the DETR-based method, all ablation studies follow the same protocol where models are trained 5 times on val set with median moderate $AP_{3D}$ reported.
@@ 4.3 Ablation Study
Segment Embeddings Threshold Val, $AP_{3D|R40}$
@@
Table 3: The design of region segmentation head. “w/o” denotes enhancing visual and depth features without segment embeddings.
Computational cost.  We analyze the computational costs of the main modules in Tab. 2 (Depth Error does not incur additional computation). Although the decoupled query and RSH modules introduce a slight increase in computational overhead, they greatly improve the performance. Our complete network achieves a runtime of 42 ms, which is comparable to recent state-of-the-art methods such as MonoCD [60] (36 ms) and OccupancyM3D [39] (112 ms).
Region segmentation head.  For the multi-scale target region probabilities predicted by RSH, we set distinct thresholds to produce segment embeddings, which serve as input tokens for the depth encoder. The results are presented in Tab. 3. The segment embeddings strength encoders’ ability to comprehend the contextual relationships among image pixels. Experiments further highlight the critical role of segment embeddings. Moreover, the threshold settings also influence 3D evaluation metrics. A threshold of 0.5 achieves superior results compared to other thresholds.
Depth Predition Mode Val, $AP_{3D|R40}$
@@
Geometric Depth ($h_{bbox}$) 9.15 7.99 5.52
Geometric Depth ($h_{c}$) 27.63 20.73 18.04
@@
Table 4: Ablation study of different depth prediction modes on KITTI val set. The weighted fusion mode integrates the results from the depth map, direct depth, and geometric depth ($h_{bbox}$). “GD” denotes the geometric depth calculated by the formula $Z_{geo}=f\times H_{3D}/h_{bbox}$.
Depth prediction mode.  Based on the same network, we compare different depth prediction modes in Tab. 4. The experiment indicates that single depth predictions, whether using direct depth, depth map, or geometric depth ($h_{bbox}$ or $h_{c}$), generally perform poorly, especially for geometric depth ($h_{bbox}$). Considering only perspective transformation on the visual surface, the moderate $AP_{3D}$ for geometric depth ($h_{bbox}$) is just 7.99% (all $AP_{3D}$ mentioned below default to the moderate level), significantly lower than other prediction modes. By fusing direct depth, depth map and geometric depth, the $AP_{3D}$ improves to 21.64%. When geometry errors are added to compensate for the projection formula, it can be observed that the performance achieves a substantial increase. The depth error wins the best performance with an $AP_{3D}$ of 22.34%. The dimension height error comes second at 22.09% due to its skewed distribution. The bounding box height error only acquires 21.18% in $AP_{3D}$, obviously inferior to other geometry errors. This discrepancy can be attributed to the non-linearity of the gradient backpropagation calculation.
Mixup3D Decoupled Query Region Segment Geometric Depth Error Val, $AP_{3D|R40}$
@@
Table 5: The main ablation study.
Overall performance.  We explore the impact of each improvement technique on the baseline. The main ablation study is shown in Tab. 5. The MonoDGP employs four key techniques: mixup3D augmentation, decoupling and initialization of 2D object queries, region enhancement with segment embeddings, and depth error prediction based on geometric depth. Compared to the baseline, the $AP_{3D}$ under three-level difficulties gain an increase by +4.20%, +2.12% and +1.80%, respectively. These experiments validate the effectiveness of each technique utilized in MonoDGP.
@@ 5 Conclusion
In this paper, we propose a novel transformer-based monocular method called MonoDGP, which adopts geometry errors to correct the projection formula. We also comprehensively analyze the properties and effectiveness of geometry errors, from the perspectives of marginal feature extraction, data distribution, and gradient propagation. Error prediction is expected to become a better alternative to multi-depth prediction, without complex multi-branch implementations. Furthermore, MonoDGP introduces a 2D visual decoder for query initialization and a region segmentation head for feature enhancement. These improvements have achieved superior performance on the KITTI benchmark. We can also extend our approach to dense depth map prediction for the target region, by only considering errors between geometric depth and surface points of the objects.
Supplementary Material
Appendix A Detailed Discussion on Depth Error
In our method, we regard the distance between the camera plane and the car’s closest wheel point as the geometric depth. However, this assumption is appropriate when the camera has the same height of the object. The height inconsistency will lead to the bias $l_{bias}$ between the actual geometric depth $Z_{geo}$ and the wheel depth $Z_{w}$. We set the height ratio $\gamma$ of the camera height $H_{cam}$ to the object height $H$ as follows:
$\gamma=\frac{H_{cam}}{H}$ (11)
We will discuss how the height ratio affects the distribution of the depth error $Z_{err}$.
Figure 7: The perspective transformation when the camera height is lower than the object height.
The vehicle is treated as a trapezoid. The closest wheel locates in the lowest position of the 2D bounding box, while the highest position in the object surface will change with the height ratio. As shown in Fig. 7, when $\gamma<1$, the wheel depth is shorter than the geometric depth, which can be expressed as:
$Z_{geo}=Z_{w}+l_{bias}$ (12)
To calculate the wheel bias, we first represent the height at the wheel point based on the similar triangle theory:
$tan\alpha=\frac{H-H_{cam}}{Z_{w}+l_{a}}=\frac{H-H_{w}}{l_{a}}$ (13)
$H_{w}=H-\frac{(H-H_{cam})\cdot l_{a}}{Z_{w}+l_{a}}$ (14)
And then we utilize $H_{w}$ to compute $l_{bias}$ :
$\frac{H}{Z_{geo}}=\frac{H_{w}}{Z_{w}}$ (15)
$l_{bias}=\frac{(H-H_{cam})\cdot Z_{w}\cdot l_{a}}{H\cdot Z_{w}+H_{cam}\cdot l_{a}}=\frac{(1-\gamma)\cdot Z_{w}\cdot l_{a}}{Z_{w}+\gamma\cdot l_{a}}$ (16)
We can express depth error as follows:
$Z_{err}=l_{b1}+l_{a}-l_{bias}=l_{b1}+\sigma_{1}\cdot l_{a}$ (17)
$\sigma_{1}=\frac{\gamma}{1-\frac{(1-\gamma)\cdot l_{a}}{Z_{w}+l_{a}}}$ (18)
where $\gamma<\sigma_{1}<1$, $l_{bias}<(1-\gamma)\cdot l_{a}$. The original depth error, which should be perspective-invariant, is calculated by the formula $Z_{err}=l_{b1}+l_{a}$. Except for the vehicle’s own attributes, $\gamma$ and $Z_{w}$ also affect the depth error. The closer $\sigma_{1}$ is to 1, the less effect it has. According to the Eq. 18, $\sigma_{1}$ reduces as $Z_{w}$ increases and $\gamma$ decreases.
Figure 8: The data distribution of the object’s central depth and dimension height on the KITTI training set.
To present the greatest impact of the height ratio, we take an extreme example based on the Fig. 8, and make $\sigma_{1}$ as smaller as possible. Specifically, we set $H_{cam}=1.5m$, $H=1.8m$, $\gamma=\frac{5}{6}$, $l_{a}=1m$, $Z_{w}=50m$. From the Eq. 16 and Eq. 18, we obtain $\sigma_{1}\approx 0.84$ and $l_{bias}\approx 0.16m$. This extreme bias value is significantly lower than the whole depth value.
Figure 9: The perspective transformation when the camera height is higher than the object height.
As shown in Fig. 9, when $\gamma>1$, the wheel depth is larger than the geometric depth, which can be expressed as:
$Z_{geo}=Z_{w}-l_{bias}$ (19)
We can achieve the wheel bias similar to the previous proof process:
$tan\beta=\frac{H_{cam}-H}{Z_{w}+l_{a}+l_{b1}+l_{b2}}=\frac{H_{w}-H}{l_{a}+l_{b1}+l_{b2}}$ (20)
$H_{w}=\frac{H\cdot Z_{w}+H_{cam}\cdot(l_{a}+l_{b1}+l_{b2})}{Z_{w}+l_{a}+l_{b1}+l_{b2}}$ (21)
$l_{bias}=\frac{(\gamma-1)\cdot Z_{w}\cdot(l_{a}+l_{b1}+l_{b2})}{Z_{w}+\gamma\cdot(l_{a}+l_{b1}+l_{b2})}$ (22)
Homogeneously, we can express depth error as follows:
$Z_{err}=l_{b1}+l_{a}+\sigma_{2}\cdot(l_{a}+l_{b1}+l_{b2})$ (23)
$\sigma_{2}=\frac{(\gamma-1)}{1+\frac{\gamma\cdot(l_{a}+l_{b1}+l_{b2})}{Z_{w}}}$ (24)
where $10<\sigma_{2}<\gamma-$ 1, $l_{bias}<(\gamma-1)\cdot(l_{a}+l_{b1}+l_{b2})$. The closer $\sigma_{2}$ is to 0, the less effect it has. According to the Eq. 24, $\sigma_{2}$ increases as $Z_{w}$ and $\gamma$ increase.
To show the height ratio’s maximum impact, we also suppose an extreme case and make $\sigma_{2}$ as larger as possible. To be more specific, we set $H_{cam}=1.5m$, $H=1.25m$, $\gamma=\frac{6}{5}$, $l_{a}+l_{b1}+l_{b2}=2m$, $Z_{w}=50m$. Based on the Eq. 22 and Eq. 24, we achieve $\sigma_{2}\approx 0.19$ and $l_{bias}\approx 0.38m$. This bias value is higher than the value calculated when $\gamma<1$, but has a slight effect on the whole depth.
In most instances, the camera height is close to the vehicle height, which means $\gamma\approx 1$ and the depth error is roughly perspective-invariant for the car category. Even if the object height is obviously different from the camera height, the network can directly learn and predict this tiny bias compared with the whole depth. The error prediction is still a simple and effective method to replace the multi-depth prediction.
Appendix B Discussion on Geometric Constraints
Previous works like Deep3DBox [36] and Shift R-CNN [37] enforce strict geometric constraints by tightly fitting projections of the 3D bounding box into the 2D box. While recent methods such as MonoGR2 [3] and GUPNet [33] formulate constraints based on geometric similarity, where under vehicle-mounted camera perspectives and fixed focal length, the object’s center depth can be uniquely determined through the proportional relationship between its 3D height and 2D projected height.
Projection-alignment constraints exhibit quadratic errors from 2D boundary localization inaccuracies, while height-ratio constraints demonstrate linear errors confined to height predictions. The former fails with truncated objects requiring full 2D contours, whereas the latter maintains functionality under partial occlusions through visible height segments. Height-ratio constraints surpass projection-alignment methods in stability (linear vs. quadratic errors), efficiency (closed-form vs. iterative), and robustness (partial vs. full contours), establishing them as core geometric priors for monocular 3D detection. Future frameworks could incorporate projection-alignment constraints as auxiliary regularizers within joint optimization.
Appendix C Detailed Loss Function
The 2D loss $L_{2D}$ adopts focal loss [28] to estimate the object categories, L1 loss to estimate the projected center $(x_{3d},y_{3d})$ and 2D sizes $(l,r,t,b)$, and GIoU loss for the bounding box. We can formulate the 2D object loss as:
$L_{2D}=\lambda_{1}L_{cls}+\lambda_{2}L_{2dsize}+\lambda_{3}L_{xy}+\lambda_{4}L_{giou}$ (25)
The 3D loss follows MonoDLE [35] to predict 3D sizes $(h_{3d},w_{3d},l_{3d})$ and orientation angle $\alpha$. As for the depth prediction, an uncertainty regression loss based on the Laplacian distribution is defined as:
${L_{depth}}=\frac{{\sqrt{2}}}{{{\sigma_{d}}}}{\left\|{\frac{{f\cdot H}}{{{h_{bbox}}}}+{Z_{err}}-{Z_{gt}}}\right\|_{1}}+\log({\sigma_{d}})$ (26)
where $\sigma_{d}$ is the standard deviation of the distribution.
We can formulate 3D object loss as:
$L_{3D}=\lambda_{5}L_{3dsize}+\lambda_{6}L_{angle}+\lambda_{7}L_{depth}$ (27)
The depth map loss $L_{dmap}$ utilizes focal loss to predict categorical foreground depth map. More detailed information about $L_{dmap}$ can be found in MonoDETR [62].
Appendix D Experiments on Other Categories
Since segment embeddings are mainly trained to distinguish between background and target, they can easily handle multiple classes without modification. Ablation studies of other categories are shown in Tab. 6. In particular, error prediction significantly improves 3D pedestrian prediction compared to cars and cyclists. This can be explained that pedestrians have consistent depth errors across orientations, whereas cyclists have irregular shapes and greater geometric uncertainty. These spatial uncertainties may degrade the effectiveness of the projection transformation.
We also compare the pedestrian and cyclist detection results in Tab. 7. Specifically, our method achieves a superior performance on all levels of difficulty for pedestrian detection, benefiting from its simple and stable geometric structures. However, the performance for the cyclist category falls short of the best.
Notably, despite these geometric challenges, our cyclist detection performance remains competitive among methods without extra training data. This underscores the generalizability of RSH module for complex categories.
Methods Val, IoU=0.5, $AP_{3D|R40}$
Pedestrian Cyclist
Easy Mod. Hard Easy Mod. Hard
MonoDGP (Ours) 13.77 10.06 7.96 12.21 6.61 5.95
w/o Segment Embeddings 13.02 9.67 7.66 10.56 5.22 4.68
w/o RSH 12.50 9.42 7.34 9.16 4.34 4.18
w/o Depth Error 9.90 7.55 6.09 11.13 5.86 5.51
Table 6: Ablation study of the pedestrian and cyclist categories on the KITTI val set.
Methods Extra data Test, IoU=0.5, $AP_{3D|R40}$
Pedestrian Cyclist
Easy Mod. Hard Easy Mod. Hard
[…]
MonoCon [30] 13.10 8.41 6.94 2.80 1.92 1.55
DEVIANT [21] 13.43 8.65 7.69 5.05 3.13 2.59
MonoDDE [26] 11.13 7.32 6.67 5.94 3.78 3.33
MonoDETR [62] 12.65 7.19 6.72 5.12 2.74 2.02
MonoDGP (Ours) None 15.04 9.89 8.38 5.28 2.82 2.65
Table 7: Comparisons of the pedestrian and cyclist categories on the KITTI test set. We bold the best results and underline the second-best results.
Appendix E Sensitivity to Initial Features
Since error prediction mode heavily relies on good geometric features, the inaccuracies of initial features can significantly impact the convergence and performance of the proposed network. The Initial features, such as 3D dimension height ($H_{3D}$) and 2D bounding box height ($h_{bbox}$), are crucial for geometric depth calculation. To analyze their individual impacts, we conduct sensitivity experiments replacing predicted $H_{3D}$ and $h_{bbox}$ with ground truth values.
As shown in Tab. 8, the perfectly accurate geometric depth improves moderate $AP_{3D}$ by up to 26.21%, highlighting the significance of these features. Compared to $h_{bbox}$, the network is more sensitive to $H_{3D}$ errors due to its inherent difficulty as a 3D property. There also exists a coupling relationship between $h_{bbox}$ and $H_{3D}$. Simultaneously replacing both features with ground truth values performs much better than replacing them individually. Current limitations mainly arise from height prediction error accumulation in the perspective projection. Improvements in monocular features, particularly for $H_{3D}$, will further enhance the performance of error prediction in the future.
Geometric Depth ( $f\times H_{3D}/h_{bbox}$ ) Val, IoU=0.7, $AP_{3D|R40}$
Ground Truth $H_{3D}$ Ground Truth $h_{bbox}$ Easy Mod. Hard
[…]
57.81 48.55 41.92
Table 8: Sensitivity study on the KITTI val set for the car category.
Depth Prediction Mode Val, IoU=0.7, $AP_{3D|R40}$
Easy Mod. Hard
Direct Depth 24.33 18.87 15.31
DAv2-small (HS) + Depth Error 11.45 9.11 7.86
DAv2-small (VK2) + Depth Error 27.04 20.48 17.52
DAv2-base (VK2) + Depth Error 27.86 21.15 18.10
Table 9: Ablation Study of pre-trained MDE. ‘DAv2’ denotes Depth Anything V2 [61] method, ‘HS’ denotes pre-trained on indoor dataset Hypersim [44], ‘VK2’ denotes pre-trianed on outdoor dataset Virtual KITTI 2 [6].
Difficulty Methods Extra $AP_{3D}$ $APH_{3D}$
All 0-30 30-50 50$\infty$ All 0-30 30-50 50$\infty$
Level_1
(IoU=0.7)
CaDDN [43] LiDAR 5.03 15.54 1.47 0.10 4.99 14.43 1.45 0.10
PatchNet [34] in [55] Depth 0.39 1.67 0.13 0.03 0.39 1.63 0.12 0.03
PCT [55] Depth 0.89 3.18 0.27 0.07 0.88 3.15 0.27 0.07
M3D-RPN [4] in [43] None 0.35 1.12 0.18 0.02 0.34 1.10 0.18 0.02
GUPNet [33] in [21] None 2.28 6.15 0.81 0.03 2.27 6.11 0.80 0.03
[…]
MonoDGP (Ours) None 4.28 10.24 1.15 0.16 4.23 10.10 1.14 0.16
Level_2
(IoU=0.7)
CaDDN [43] LiDAR 4.49 14.50 1.42 0.09 4.45 14.38 1.41 0.09
PatchNet [34] in [55] Depth 0.38 1.67 0.13 0.03 0.36 1.63 0.11 0.03
PCT [55] Depth 0.66 3.18 0.27 0.07 0.66 3.15 0.26 0.07
M3D-RPN [4] in [43] None 0.35 1.12 0.18 0.02 0.33 1.10 0.17 0.02
GUPNet [33] in [21] None 2.14 6.13 0.78 0.02 2.12 6.08 0.77 0.02
[…]
MonoDGP (Ours) None 4.00 10.20 1.13 0.15 3.96 10.08 1.12 0.15
Level_1
(IoU=0.5)
CaDDN [43] LiDAR 17.54 45.00 9.24 0.64 17.31 44.46 9.11 0.62
PatchNet [34] in [55] Depth 2.92 10.03 1.09 0.23 2.74 9.75 0.96 0.18
PCT [55] Depth 4.20 14.70 1.78 0.39 4.15 14.54 1.75 0.39
M3D-RPN [4] in [43] None 3.79 11.14 2.16 0.26 3.63 10.70 2.09 0.21
GUPNet [33] in [21] None 10.02 24.78 4.84 0.22 9.94 24.59 4.78 0.22
[…]
MonoDGP (Ours) None 12.36 31.12 5.78 1.24 12.18 30.68 5.71 1.22
Level_2
(IoU=0.5)
CaDDN [43] LiDAR 16.51 44.87 8.99 0.58 16.28 44.33 8.86 0.55
PatchNet [34] in [55] Depth 2.42 10.01 1.07 0.22 2.28 9.73 0.97 0.16
PCT [55] Depth 4.03 14.67 1.74 0.36 4.15 14.51 1.71 0.35
M3D-RPN [4] in [43] None 3.61 11.12 2.12 0.24 3.46 10.67 2.04 0.20
GUPNet [33] in [21] None 9.39 24.69 4.67 0.19 9.31 24.50 4.62 0.19
[…]
MonoDGP (Ours) None 11.71 31.02 5.61 1.17 11.56 30.58 5.54 1.15
[…]
Appendix F Initial Depth from Pre-trained MDE
Monocular depth estimation (MDE) models have developed for many years. We can also utilize the pre-trained MDE to provide a roughly approximate surface depth, similar to geometric depth, which may render the learning problem even simpler.
To explore this possibility, we exploit Depth Anything V2 [61] to generate depth maps. Based on the initial metric depth, error prediction can achieve better performance compared to direct prediction in Tab. 6. However, MDE heavily relies on pre-trained datasets, while geometric depth relies on its own attributes without additional parameters. This will limit the generalization of achieving initial depth from pre-trained MDE.
Appendix G Experiments on Waymo Open Dataset
Waymo [52] evaluates objects at Level_1 and Level_2, which are determined by the number of LiDAR points within their 3D bounding boxes. The experiments is conducted across three distance ranges: [0, 30), [30, 50), and [50, $\infty$) meters. Performance on the Waymo dataset is assessed by average precision $AP_{3D}$ and average precision weighted by heading $APH_{3D}$.
We follow the DEVIANT [21] split to generate 52,386 training and 39,848 validation images by sampling every third frame. For fairness, we mainly compare with methods using the same split in Tab. 10. Our method achieves state-of-the-art performance without extra data across all ranges, particularly for distant objects. These results further validate the effectiveness and generalizability of MonoDGP. It is worth noting that CaDDN [43]’s performance is better than MonoDGP, this discrepancy may be attributed to different dataset splits and introduction of LiDAR data.
Appendix H Qualitative Discussion and Visualization
To provide a more intuitive comparison between our method and the baseline models, we visualize some 3D detection results from both the camera view and the bird’s-eye view on the KITTI validation set. As shown in Fig. 10, our method demonstrates superior performance on distant and length-occluded objects.
However, since error prediction is affected by the initial accuracy of geometric depth, which is calculated from height relationships, height occlusion remains a challenge for our method. For the leftmost vehicle in the third example of Fig. 10, bushes block out its lower part, weakening the accuracy of height and consequently propagating errors to depth prediction. This failure case highlights the need for further improvements in handling height occlusion, potentially through the integration of additional contextual information or more robust occlusion-aware models.
Figure 10: Qualitative results on KITTI validation set. (a) MonoCD (b) MonoDETR (c) MonoDGP (ours). In each group of images, the first row shows the camera view, and the second row shows the bird’s-eye view. Green represents the ground truth of boxes, while Red represents the prediction results. We also circle some objects to highlight the difference between the baseline model and our method.